Thomas collected 29 responses to a survey and used a two-tailed test to establish a significance level of 0.20 and calculate a 90% confidence interval.

Using the t-table (located in the tutorial), what is the critical t-value used for calculating a 90% confidence interval with 28 degrees of freedom?
Answer #1

Sample size = 29

Degree of freedom = 29-1 = 28

Level of significance =

t critical value is =1.701.................using t table or by using Excel command =TINV(0.1,28)